OpenStudy (anonymous):

Simplify: the square root of 300x^10 A. 10x^8 and the square root of 3 B. 100x^8 and the square root of 3 C. 150x^8 D. 25x^4 and the square root of 12

OpenStudy (anonymous):

none of these are correct, could the question be 300x^16

OpenStudy (anonymous):

If so the answer is A

OpenStudy (anonymous):

The equation is: \[\sqrt{300x ^{10}}\]

OpenStudy (anonymous):

So is A the answer then?

OpenStudy (anonymous):

If it is a 10, none of them is the answer

OpenStudy (anonymous):

How does that work out then? It has to be one of them.. because there is no choice you can pick "none of the above"

OpenStudy (anonymous):

typo? Ask the teacher or computer program about their typo. Is there a complaints department??

OpenStudy (anonymous):

There is no typo... I will re-write the entire thing in equation form.. hold on, please.

OpenStudy (anonymous):

Simplify: \[\sqrt{300x ^{10}}\] A. \[10x ^{8}\sqrt{3}\] B. \[100x ^{8}\sqrt{3}\] C. \[150x ^{8}\] D. \[25x ^{4}\sqrt{12}\]

OpenStudy (anonymous):

Are those 5's and not 8's? If so, then A Is there a font complaint department?

OpenStudy (anonymous):

Okay clearly, ONE of these answers HAS to be correct in some way, those are the ONLY choices they give. And there is no "font complaint department", etc. Three of those are 8's, and one is a 4, in the answer choices. Get glasses if you can't tell the difference.

OpenStudy (anonymous):

If there eights then none of them are correct. I used grandma's glasses!

OpenStudy (anonymous):

Your teacher/the computer program can make mistakes. My math is correct and my eyes work

OpenStudy (anonymous):

Then I guess I'll just have to guess an answer huh?

OpenStudy (anonymous):

or ask your teacher ... I believe those are 5's and the answer is A

OpenStudy (anonymous):

It says they are 8's on the worksheet...

OpenStudy (anonymous):

but its whatever , im done with this conversation .

OpenStudy (anonymous):

I believe with the typo the answer is A

OpenStudy (anonymous):

well thank you

OpenStudy (anonymous):

we tried ...